you purchase 100 shares of Mun Tec ltd on a 55% margin when the shares are selling at K20 each. The Lusaka Stick Exchange broker charges you 10% annual interest and commissions are 3% of the total stock value on both the purchase and sale. If a year later you receive a K0.5 per share dividend and sell the stock for K27. What is your rate of return on this investment?
1
Expert's answer
2021-09-17T09:20:29-0400

Solution:

Return On Capital (ROI) = "\frac{Net \; return\; on\; Investment}{Cost\; of\; Investment}\times 100"

Net return on Investment = Sales value + Dividends – Annual interest – Commissions

Sales value = K27"\times"100 shares = 2,700

Purchase/Cost of investment = K20"\times"100 shares = 2,000

Dividends = K0.5"\times"100 shares = 50


Annual Interest = (10"\% \times"2,000) + (10"\% \times"2,700) = 200 + 270 = 470


Commissions = (3"\% \times"2,000) + (3"\% \times" 2,700) = 60 + 81 = 141


Net return on Investment = (2,700 + 50) – (2,000 – 470 – 141) = 139


Return On Capital (ROI) = "\frac{139}{2000} \times 100\% = 6.95\%"